OpenStudy (anonymous):

What is r(remainder) when (4x3 + 2x2 – 18x + 38) ÷ (x – 3)

OpenStudy (anonymous):

Do you know how to do long division of polynomials? or synthetic division? either of those two methods will give you the answer.

OpenStudy (anonymous):

i got 110

OpenStudy (anonymous):

lets see: \[4(3)^3+2(3)^2-18(3)+38 = 110\]so yes you are correct :)

OpenStudy (anonymous):

What is the solution to the equation log4 x + log4 2 = 3 ? do u know the answer to this?

OpenStudy (anonymous):

hmm, you can use the properties of logs to combine the two being added. Use:\[\log_b(x)+\log_b(y)=\log_b(xy)\]

OpenStudy (anonymous):

That would give you:\[\log_4(2x)=3\]

OpenStudy (anonymous):

okay what did you get as an answer

OpenStudy (anonymous):

32 i believe, since:\[\log_4(2x)=3\Longrightarrow 4^3 = 2x\iff 64=2x\iff x = 32\]

OpenStudy (anonymous):

i got 32 also. how about this? If f(x) = 5x2 – 8x + 13 and g(x) = 6x – 5, what is f(x) + g(x)?